The length of a rectangular garden is 3 yd more than twice its width. The perimeter of the garden is 36 yd, what are the width and length of the garden?
Answer by rfer(16322)

l=2x+3=13
w=x=5
---------------
2(2x+3)+2x=36
4x+6+2x=36
6x=30
x=5